ABC, Inc. purchased an equipment at time=0 for $140,132. The shipping and installation costs were $15,535. The equipment is classified as a 7-year MACRS property. The investment in net working capital at time=0 was $14,402 which would be recouped at the end of the project. The project life is six years. At the end of the sixth year, the company will sell the equipment for $6,548. The annual cash flows are $95,121. What is the cash flow of the project in Year 6? That is solve for CF6. Assume that the tax rate is 22%. The MACRS allowance percentages are as follows, starting with Year 1: 14.29, 24.49, 17.49, 12.49, 8.93, 8.92, 8.93, and 4.46 percent. Note: In the last year of the project, the Total Cash Flow = Operating Cash Flow + Terminal Cash Flow Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

Operating cash flow in Year 6 = 95121

Written down value of Equipment = (140,132+15535)*13.39% = $20,843.8113

Selling price = $6548

Loss on Sale = $14,295.8113

Tax savings = $3145.08

Hence, salvage value after tax = 6548+3145.08 = $9693.08

Recovery of working capital = $14402

Terminal cash flow = $24,095.08

Cash flow in Year 6 = 119,216.08